Solve the system:
x^2+y^2=63
x^2-3y^2=27

Answer by Alan3354(69443)

x^2+y^2=63
x^2-3y^2=27
----------- Subtract
4y^2 = 36
y = +/- 3
----------
Sub for y and find values of x
x^2 + 9 = 63
x = +/- 3sqrt(2)
Use the + and minus values to get the 4 points of intersection.